For an isotropic solid, show that the coefficients of expansion α, 𝛽 and 𝛾 are related as

(a) 𝛽 = 2𝛼

(b) 𝛾 = 3?


1
Expert's answer
2021-10-06T08:12:39-0400

We know that linear expansion

"l_1 = l(1+\alpha \Delta T)...(i)"

Area expansion

"A_1=A(1+\beta \Delta T)=l^2(1+\beta \Delta T)...(ii)"

Volumetric expansion

"V_1= V(1+\gamma \Delta T)...(iii)"

But we know that area "(A_1)=l_{length}\times l_{base}=l^2(1+\alpha \Delta T)^2"

Now, taking the bi-nominal expansion,

"A_1=l^2(1+2\alpha)...(iv)"

from equation (ii) and (iv)

So, "\beta = 2\alpha"

Similarly,

"V_1= A_1\times l_1=l^2(1+3\alpha)...(v)"

Hence, from equation (i) and (v)

So, "\gamma = 3\alpha"